Cost of Materials Issuances Under the FIFO Method An incomplete subsidiary ledger of materials inventory for May is as follows:

Complete the materials issuances and balances for the materials subsidiary ledger under FIFO.

Received Issued Balance
Receiving
Report
Number
Quantity Unit
Price
Materials
Requisition
Number
Quantity Amount Date Quantity Unit
price
Amount
May 1 310 $6 $1,860
26 220 $8 May 4
108 350 $ May 10
32 150 10 May 21
121 210 May 27
Answer #1
Received Issued Balance
Receiving Quantity Unit Materials Quantity Amount Date Quantity Unit Amount
Report Price Requisition price
Number Number
1-May 310 $6 $1,860
26 220 $8 4-May 310 $6 $1,860
220 $8 $1,760
108 350 2180 10-May 180 $8 $1,440
32 150 10 21-May 180 $8 $1,440
150 $10 $1,500
121 210 1740 27-May 120 $10 $1,200
Workings:
Issued:
108 = (310*6)+(350-310)*8 = $2180
121 = (180*8)+(210-180)*10 = $1740